Please help.
The difference of 2 numbers is 6. The second is 9 less than 2 times the first. What are the 2 numbers?

Answer by jayanthihemkos@yahoo.com(87)

The difference of 2 numbers is 6. The second is 9 less than 2 times the first. What are the 2 numbers?

Let the first number = x
The second number = 2x-9
The difference of the two number = 6

x-(2x-9) = 6
x - 2x +9 = 6
-x +9 = 6
-x +9 -9 = 6-9
-x = -3
x = 3 (-ve sign on bothside gets cancelled)

CHECK
x-(2x-9) = 6
3-[2(3)-9] = 6
3 - [6-9] = 6
3-(-3) = 6
3+3 = 6
6= 6
LHS = RHS
Hence x = 3 is the solution